OpenStudy (anonymous):

4x − 2y − z, if x = 2, y = 3, and z = −10

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

Evaluate the following expression using the values given.

OpenStudy (anonymous):

So just plug in your values for the letters. x=2 so wherever you see an x put a 2 instead y=3 so wherever you see a y put a 3 instead z=-10 so wherever you see a z put a -10 instead Then just do basic math. 4x-2y-z = 4(2) - 2(3) - (-10) = 8 - 6 + 10 = 8 - 16 = -8 Remember that a double negative is a positive, so -(-10) = 10. And remember PEMDAS for which order to do your math: Parentheses, Exponents, Multiplication, Division, Addition, Subtraction. You had 8 - 6 + 10 so you Add the 6 + 10 first before Subtracting that from the 8. It's been a while since I've had to break out the word PEMDAS but I believe that's correct.